# Camouflaged Object Tracking: A Benchmark

### v2

@@ I Introduction
In recent years, the field of visual tracking has made significant advances, largely due to the widespread use of large-scale training datasets. The richness of these datasets has facilitated the development of high-precision and robust tracking algorithms [1, 2, 3, 4, 5, 6, 7, 8]. Although substantial progress has been made in general object tracking, research into more specialized and challenging scenarios remains limited. Among these, tracking camouflaged objects is particularly difficult because these targets seamlessly blend into their surroundings, making them hard to detect with the naked eye and significantly increasing tracking complexity [9, 10].
@@
Distinguishing camouflaged objects from their backgrounds in these applications is crucial, but it involves significant challenges like background clutter, boundary blurring, texture mimicry, color similarity, varying lighting conditions, and occlusion [21, 9, 10]. Addressing these challenges through innovative approaches and advanced technologies can lead to significant advancements in applications involving camouflaged objects [22, 23, 24, 25, 26]. In recent years, numerous computer vision techniques have been developed to address the challenges posed by camouflaged objects [27, 26]. One prominent direction is Camouflaged Object Detection (COD), which identifies objects blending into their surroundings due to similar textures, shapes, or colors. COD locates these hidden objects within images or videos [28, 29]. Camouflaged Object Segmentation (COS) takes this further by not only detecting but also accurately separating these objects from their backgrounds. COS involves classifying each pixel to determine if it belongs to the object or the background, offering detailed information about the object’s shape and edges [30]. The ongoing research and development in COD and COS continue to push the boundaries, promising further innovations and applications that tackle complex real-world challenges related to camouflaged objects. The ongoing research in COD and COS continue to push the boundaries, promising further innovations and applications that tackle real-world challenges related to camouflaged objects.
@@ II-A Visual Tracking Algorithms
DCF-based trackers treat the visual tracking problem as an online regression task, leveraging the Parseval theorem and Fast Fourier Transform (FFT) for efficient computation. The MOSSE (Minimum Output Sum of Squared Error) filter [35] is an early representative in this domain. Since its introduction, significant advancements have been made. For example, [36, 37] introduced additional scale filters to address variations in object scale, while [38, 39] employed regularization techniques to enhance tracker robustness. Furthermore, extensions of the DCF framework, such as translation-invariant techniques, continue to evolve [40]. However, due to the limited representational capacity of handcrafted features, DCF-based trackers still face challenges in robustness under complex conditions.
To overcome the limitations of DCF-based trackers, deep learning-based trackers have garnered significant attention in recent years. These trackers utilize neural networks to automatically learn features, substantially improving tracking performance. SiamFC [41] pioneered this field by applying Siamese networks [42] to visual tracking. SiamFC learns a nonlinear similarity function through offline training and performs correlation operations on candidate regions and the initial object state to generate response maps for the search area. Building on this, the SiamRPN algorithm introduced a candidate region generation network structure with classification and regression branches to enhance tracking speed. Subsequent research, such as SiamRPN++ [43], SiamR-CNN [44], SiamCAR [45], and Ocean [46], has further improved the accuracy and robustness of Siamese trackers by incorporating techniques such as attention mechanisms, backbone network optimization, and enhanced contextual information.
@@ II-B Visual Tracking Benchmarks
Specific benchmarks: Specific benchmark datasets aim to evaluate tracking performance for particular types of objects. For example, UAV123 [65] focuses on low-altitude drone object tracking, containing 123 sequences captured by drones. CDTB [66] and PTB [67] are designed to evaluate tracking performance in RGB-D videos, where D represents depth images. TOTB [52] specializes in tracking transparent objects, collecting 225 videos from 15 categories of transparent objects. Recently, TSFMO [68] has been introduced for tracking small and fast-moving objects, while TRO [69] focuses on tracking reflected objects.
@@ II-C Dealing with camouflaged Objects in Vision
Camouflaged objects are difficult to distinguish from their backgrounds because they blend into their surroundings. Unlike regular objects, they have similar textures, colors, and patterns as their environment, creating visual ambiguity. Recently, many computer vision techniques have been developed to tackle these challenges. For instance, Fan et al. [28] introduced the task of Camouflaged Object Detection (COD) and established a dataset COD10K with 10,000 images, along with a simple yet effective framework SINet for efficient and precise detection of camouflaged objects. Liu et al. [29] proposed a scale-adaptive correlation filter algorithm that integrates FHOG and LBP features for Camouflaged Object Segmentation (COS). Cheng et al. [22] developed a Video Camouflaged Object Detection (VCOD) framework that captures inter-frame motion and uses a spatio-temporal transformer for consistency, along with the MoCA-Mask dataset providing detailed pixel-level annotations. Hu et al. [70] tackled visual camouflage challenges in UAV surveillance with a novel trajectory planning control framework. Despite these advancements, Camouflaged Object Tracking (COT), which focuses on consistently identifying camouflaged objects in videos while adapting to changes in position, scale, and appearance, remains under-explored. The absence of publicly available datasets for camouflaged object tracking has further constrained the progress of research and algorithm development in this field. To bridge this gap, we introduce the Camouflaged Object Tracking Dataset (COTD), specifically designed to support the development and evaluation of algorithms tailored for tracking camouflaged objects.
@@ III-A Data Collection
Benchmark Videos Total frames Object classes Attr. Resolution Feature
OTB100 100 81K 16 11 128$\times$96 $\sim$ 800$\times$336 short-term
LaSOT* 280 685K 85 14 202$\times$360 $\sim$ 1280$\times$720 category balance
GOT-10k* 420 56K 84 6 270$\times$480 $\sim$ 3840$\times$2160 generic
VOT 62 20K 37 9 320$\times$180 $\sim$ 1280$\times$720 annual
COTD 200 80K 29 12 244$\times$578 $\sim$ 2400$\times$1080 camouflaged
By contrast, our proposed COTD is the first benchmark dataset to focus on tracking camouflaged objects. A summarized comparison with existing benchmarks is reported in Table II.
@@ IV Method
IV-A Multi-Level Feature Sensing Encoder
@@
IV-B Overall Architecture
@@ V-B Evaluation Results
Attention mechanism analysis: To make sense of the advantage of the propose method, we explore the allocation of attention in the task of Camouflaged Object Tracking and compare the performance of the improved HIPTrack-MLS model with the original HIPTrack model. Fig. 12 intuitively displays three types of visual representations: the original image, the attention map of the original HIPTrack model, and the attention map of the HIPTrack-MLS model. The original images in the first row demonstrate the camouflage of the object in a complex background, which poses a significant challenge to tracking algorithms. The high degree of visual integration between the object and the background increases the difficulty of tracking. The second row shows the attention maps generated by the original HIPTrack model. These maps reveal some limitations of the original model when handling camouflaged objects, especially when the similarity between the object and the background is high, the original model may struggle to maintain precise tracking of the object. The third row, however, showcases the attention maps of the HIPTrack-MLS model, which clearly highlight the target object. Even when the object is camouflaged, the model can accurately identify and focus attention on the key features of the object. This improved attention mechanism significantly enhances tracking accuracy, especially in complex environments.
@@
Tracker LaSOT TrackingNet COTD
@@
Performance on generic datasets: To validate whether HIPTrack-MLS’s improvements are specifically beneficial for camouflaged object tracking or represent a general enhancement in tracking capability, we conduct comprehensive evaluations on standard tracking benchmarks including LaSOT and TrackingNet, alongside our COTD dataset. The results are presented in Table VI. On the LaSOT dataset, HIPTrack-MLS achieves comparable performance to the baseline HIPTrack, with precision and success scores of 0.793 and 0.724, respectively, showing a slight marginal difference compared to HIPTrack’s 0.795 and 0.727. Similarly, on TrackingNet, HIPTrack-MLS demonstrates minimal improvement with precision and success scores of 0.837 and 0.845, compared to HIPTrack’s 0.838 and 0.845. However, on the COTD dataset, HIPTrack-MLS shows a significant performance gain, improving the success rate from 0.680 to 0.692 and precision from 0.515 to 0.525. This notable improvement on COTD, coupled with maintained performance on general tracking benchmarks, suggests that our multi-level feature sensing architecture specifically addresses the challenges of camouflaged object tracking without compromising general tracking capability.

### v3

@@ I Introduction
In recent years, the field of visual tracking has made significant advances, largely due to the widespread use of large-scale training datasets. The richness of these datasets has facilitated the development of high-precision and robust tracking algorithms [1, 2, 3, 4, 5, 6, 7, 8]. Although substantial progress has been made in general object tracking, research into more specialized and challenging scenarios remains limited. Among these, tracking camouflaged objects is particularly difficult because these targets seamlessly blend into their surroundings, making them hard to detect with the naked eye and significantly increasing tracking complexity [9, 10]. While existing tracking methods have achieved impressive results on regular objects through various techniques such as high-level semantics [11], block-based correlation filters [12], spatio-temporal feature learning [13], and long-term tracking frameworks [14], their performance degrades significantly when dealing with camouflaged targets due to the inherent visual ambiguity between the target and background.
@@
Distinguishing camouflaged objects from their backgrounds in these applications is crucial, but it involves significant challenges like background clutter, boundary blurring, texture mimicry, color similarity, varying lighting conditions, and occlusion [25, 9, 10]. Addressing these challenges through innovative approaches and advanced technologies can lead to significant advancements in applications involving camouflaged objects [26, 27, 28, 29, 30]. In recent years, numerous computer vision techniques have been developed to address the challenges posed by camouflaged objects [31, 30]. One prominent direction is Camouflaged Object Detection (COD), which identifies objects blending into their surroundings due to similar textures, shapes, or colors. COD locates these hidden objects within images or videos [32, 33]. Camouflaged Object Segmentation (COS) takes this further by not only detecting but also accurately separating these objects from their backgrounds. COS involves classifying each pixel to determine if it belongs to the object or the background, offering detailed information about the object’s shape and edges [34]. The ongoing research in COD and COS continue to push the boundaries, promising further innovations and applications that tackle real-world challenges related to camouflaged objects.
@@ II-A Visual Tracking Algorithms
DCF-based trackers treat the visual tracking problem as an online regression task, leveraging the Parseval theorem and Fast Fourier Transform (FFT) for efficient computation. The MOSSE (Minimum Output Sum of Squared Error) filter [39] is an early representative in this domain. Since its introduction, significant advancements have been made. For example, [40, 41] introduced additional scale filters to address variations in object scale, while [42] and [43] employed regularization techniques to enhance tracker robustness. Furthermore, extensions of the DCF framework, such as translation-invariant techniques, continue to evolve [44]. However, due to the limited representational capacity of handcrafted features, DCF-based trackers still face challenges in robustness under complex conditions.
To overcome the limitations of DCF-based trackers, deep learning-based trackers have garnered significant attention in recent years. These trackers utilize neural networks to automatically learn features, substantially improving tracking performance. SiamFC [45] pioneered this field by applying Siamese networks [46] to visual tracking. SiamFC learns a nonlinear similarity function through offline training and performs correlation operations on candidate regions and the initial object state to generate response maps for the search area. Building on this, the SiamRPN algorithm introduced a candidate region generation network structure with classification and regression branches to enhance tracking speed. Subsequent research, including SiamRPN++ [47], SiamR-CNN [48], SiamCAR [49], and Ocean [50], has enhanced the accuracy and robustness of Siamese trackers through techniques like attention mechanisms in SiamCAR, backbone optimization in SiamRPN++, ResNet-50 feature enhancement in Ocean, and two-stage detection in SiamR-CNN.
@@ II-B Visual Tracking Benchmarks
Specific benchmarks: Specific benchmark datasets aim to evaluate tracking performance for particular types of objects. For example, UAV123 [69] focuses on low-altitude drone object tracking, containing 123 sequences captured by drones. CDTB [70] and PTB [71] are designed to evaluate tracking performance in RGB-D videos, where D represents depth images. TOTB [56] specializes in tracking transparent objects, collecting 225 videos from 15 categories of transparent objects. Recently, TSFMO [72] has been introduced for tracking small and fast-moving objects, while TRO [73] focuses on tracking reflected objects.Additionally, BihOT [74] contains 49 videos and 42K frames, focusing on hyperspectral camouflage. UW-COT220 [75] focuses on underwater camouflage
@@ II-C Dealing with camouflaged Objects in Vision
Camouflaged objects are difficult to distinguish from their backgrounds because they blend into their surroundings. Unlike regular objects, they have similar textures, colors, and patterns as their environment, creating visual ambiguity. Recently, many computer vision techniques have been developed to tackle these challenges. For instance, Fan et al. [32] introduced the task of Camouflaged Object Detection (COD) and established a dataset COD10K with 10,000 images, along with a simple yet effective framework SINet for efficient and precise detection of camouflaged objects. Liu et al. [33] proposed a scale-adaptive correlation filter algorithm that integrates FHOG and LBP features for Camouflaged Object Segmentation (COS). Yue et al. [76] and Chen et al. [77] have further refined segmentation accuracy by integrating multi-scale feature fusion techniques. Cheng et al. [26] developed a Video Camouflaged Object Detection (VCOD) framework that captures inter-frame motion and uses a spatio-temporal transformer for consistency, along with the MoCA-Mask dataset providing detailed pixel-level annotations. Additional research by Wang et al. [78] have demonstrated the potential of leveraging discriminative feature learning for enhancing detection robustness in complex environments. Despite these advancements, Camouflaged Object Tracking (COT), which focuses on consistently identifying camouflaged objects in videos while adapting to changes in position, scale, and appearance, remains under-explored. The absence of publicly available datasets for camouflaged object tracking has further constrained the progress of research and algorithm development in this field. To bridge this gap, we introduce the Camouflaged Object Tracking Dataset (COTD), specifically designed to support the development and evaluation of algorithms tailored for tracking camouflaged objects.
@@ III-A Data Collection
Compared to existing benchmark datasets, the COTD (ours) dataset focuses on tracking camouflaged objects, which presents a unique challenge in the field of visual single-object tracking. A summarized comparison with existing benchmarks is reported in Table II. The COTD dataset includes 200 videos and 80,000 frames, surpassing in quantity the 100 videos and 59K frames of OTB100 [79], as well as the 70 videos and 15.8K frames of DTB70 [80]. Furthermore, compared to BihOT [74] and UW-COT220 [75], COTD has a higher focus on camouflage tracking. BihOT contains 49 videos and 42K frames, specializing in hyperspectral camouflage, while UW-COT220, although it includes 220 videos and 159K frames, focuses on underwater camouflage. However, the scale and diversity of the COTD dataset enable it to better simulate the complexity and challenges of tracking camouflaged objects in real-world scenarios. COTD specializes in tracking camouflaged objects in various environments, contrasting with the application scenarios of BihOT and UW-COT220.
@@
Benchmark Year Videos Total frames classes Attr. Avg.frames Resolution Visual annotation Feature
OTB100 [79] 2015 100 59K 16 11 590 128$\times$96 $\sim$ 800$\times$336 BBox short-term
UAV123 [88] 2016 123 113K 9 12 915 720$\times$480 $\sim$ 1280$\times$720 BBox UAV
DTB70 [80] 2017 70 15.8K - 11 225 1280$\times$720 $\sim$ 1280$\times$720 BBox UAV
UAVDT [89] 2018 50 37.1K 3 8 742 960$\times$540 $\sim$ 1024$\times$540 BBox UAV
UAVTrack112 [90] 2021 112 100K 32 13 895.65 1280$\times$720 $\sim$ 1920$\times$1080 BBox UAV
WebUAV3M [91] 2022 4500 3.3M 223 17 710 608$\times$320 $\sim$ 1982$\times$1088 BBox UAV
TrackingNet* [66] 2018 511 226K 27 15 441.47 270$\times$360 $\sim$ 1280$\times$720 BBox large scale
LaSOT* [68] 2019 280 685K 85 14 2447.71 202$\times$360 $\sim$ 1280$\times$720 BBox category balance
GOT-10k* [67] 2019 180 23K 84 6 133.33 270$\times$480 $\sim$ 3840$\times$2160 BBox generic
VOT [92] 2016 60 21K 37 9 357.58 320$\times$180 $\sim$ 1280$\times$720 BBox annual
TSFMO250 [72] 2022 250 51K 26 12 202.72 720$\times$404 $\sim$ 1310$\times$720 BBox small and fast moving
TRO [73] 2024 200 70K - 12 349.03 244$\times$578 $\sim$ 2400$\times$1080 BBox reflected
LLOL [93] 2024 269 133K 32 12 493.66 544$\times$960 $\sim$ 3840$\times$2160 BBox low-light
MoCA-Mask [26] 2022 87 23K 45 - 263.67 480$\times$360 $\sim$ 1920$\times$1080 Mask camouflaged
BihoT [74] 2024 49 42K - 9 855.35 - BBox hyperspectral camouflaged
UW-COT220 [75] 2025 220 159K 96 - 722 320$\times$240 $\sim$ 3840$\times$2160 BBox+Mask underwater camouflaged
COTD (Ours) 2025 200 80K 20 12 406.55 270$\times$480 $\sim$ 2488$\times$1152 BBox camouflaged
@@ IV Method
IV-A Overall Architecture
@@
IV-B Multi-Level Feature Sensing Encoder
@@ V-B Evaluation Results
Tracker LaSOT [98] TrackingNet [99] COTD
@@
Performance on generic datasets: To validate whether HIPTrack-MLS’s improvements are specifically beneficial for camouflaged object tracking or represent a general enhancement in tracking capability, we conduct comprehensive evaluations on standard tracking benchmarks including LaSOT [98] and TrackingNet [99], alongside our COTD dataset. The results are presented in Table VI. On the LaSOT dataset, HIPTrack-MLS achieves comparable performance to the baseline HIPTrack, with precision and success scores of 0.793 and 0.726, respectively, showing a slight marginal difference compared to HIPTrack’s 0.795 and 0.727. Similarly, on TrackingNet, HIPTrack-MLS demonstrates minimal improvement with precision and success scores of 0.837 and 0.845, compared to HIPTrack’s 0.838 and 0.845. However, on the COTD dataset, HIPTrack-MLS shows a significant performance gain, improving the success rate from 0.680 to 0.692 and precision from 0.515 to 0.525. This notable improvement on COTD, coupled with maintained performance on general tracking benchmarks, suggests that our multi-level feature sensing architecture specifically addresses the challenges of camouflaged object tracking without compromising general tracking capability.
@@
V-C Ablation study
Attention mechanism analysis. To make sense of the advantage of the propose method, we explore the allocation of attention in the task of Camouflaged Object Tracking and compare the performance of the improved HIPTrack-MLS model with the original HIPTrack model. Fig. 12 intuitively displays three types of visual representations: the original image, the attention map of the original HIPTrack model, and the attention map of the HIPTrack-MLS model. The original images in the first row demonstrate the camouflage of the object in a complex background, which poses a significant challenge to tracking algorithms. The high degree of visual integration between the object and the background increases the difficulty of tracking. The second row shows the attention maps generated by the original HIPTrack model. These maps reveal some limitations of the original model when handling camouflaged objects, especially when the similarity between the object and the background is high, the original model may struggle to maintain precise tracking of the object. The third row, however, showcases the attention maps of the HIPTrack-MLS model, which clearly highlight the target object. Even when the object is camouflaged, the model can accurately identify and focus attention on the key features of the object. This improved attention mechanism significantly enhances tracking accuracy, especially in complex environments.